Grade 12General Physics

determine the magnitude and direction of the resultant of the forces of magnitude 12N and 9N acting at a point. if the angle between two forces is 30 degrees

mafuta calvine
7 Years agoGrade 12

1 Answer

Arun
7 Years ago
Dear student
 
Resultant = R (let)
 
R^2 = 144 + 81 + 2*12 * 9 * sin30
 
R^2 = 333
 
R = sqrt(333) N
 
for direction
 
tan theta = 9 sin30 / (12 + 9 cos e0)